(2a+3b)+4c=2a+(3b+4c)
Found 2 solutions by stanbon, Edwin McCravy:
Answer by stanbon(75887)

(2a+3b)+4c=2a+(3b+4c)
-----
Subtract 2a+3b+4c from both sides and you get:
0 = 0
----
The Law illustrated by your problem is
the Associative Law of addition.
===================================
Cheers,
Stan H.

Answer by Edwin McCravy(20055)


 (2a+3b)+4c=2a+(3b+4c)

That's called the "associative principle" of 
addition because on the left side, the parentheses 
associates the first two terms, 2a and 3b, together, 
whereas on the right side the parentheses associates 
the last two terms, the 3b and the 4c, together.

The associative principle says that when there are
three terms to add, it doesn't matter which two you 
associate together with parentheses, the answer will 
always come out the same if you substituted numbers 
for all those letters.

Suppose a stood for 3, b stood for 4, and c stood for 2

Then the left side of  (2a+3b)+4c=2a+(3b+4c) would be

 (2a+3b)+4c

and after substituting we'd have:

 (2*3+3*4)+4*2
 ( 6 + 12)+ 8
    18    + 8
        26

and the right side of  (2a+3b)+4c=2a+(3b+4c) would be

 2a+(3b+4c)

and after substituting we'd have:

 2*3+(3*4+4*2)
  6 +( 12+ 8 )
  6 +   20
     26

The work was different because we added different
numbers, but the final answer came out 26 in both.  
That demonstrates why if we have three
terms to add together it doesn't matter
whether we associate the first two 
or the last two.  Only the work will
be different, but not the final answer.

Edwin